OpenStudy (anonymous):

medal to answer ! what is the value of 6*8-4[7+(8-3)]?

jagr2713 (jagr2713):

Simplify the following: 6×8-4 (7+8-3) Evaluate 7+8. 7+8 = 15: 6×8-4 15-3 Subtract 3 from 15. which is 12×8-4×12 Multiply 6 and 8 together. 6×8 = 48: 48-4×12 Multiply -4 and 12 together. -4×12 = -48: 48+-48 Look for two terms that sum to zero. 48-48 = 0

OpenStudy (anonymous):

the answer is 0 ? @jagr2713

jagr2713 (jagr2713):

yes